OpenStudy (anonymous):

Find the upper-bound for the summation, such that: (using the lowest possible integer)

OpenStudy (amistre64):

it feels like there is something missing in the post

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\[\sum_{n=4200}^{x}n ^{1.5} \ge 1.0 \times 10^{8}\]

OpenStudy (anonymous):

not sure if possible. I could solve it if it was something like 5n

OpenStudy (amistre64):

does this have anything to do with a remainder thrm?

OpenStudy (anonymous):

I want to solve it for a game. Putting in random numbers, the closest integer I could get to have the same as close as possible to 100 million is 4563

OpenStudy (anonymous):

*starting at 4220 >_>

OpenStudy (anonymous):

Wolfram gives me weird zeta functions or harmonic numbers

OpenStudy (amistre64):

10^8, or 10^7?

OpenStudy (anonymous):

100 million, so 10^8 (if I did that right)

OpenStudy (amistre64):

\[\int_{4200}^{x}x^{1.5}dx=x^{2.5}/2.5-(4200)^{2.5}/2.5\] \[x^{2.5}/2.5-(4200)^{2.5}/2.5=10^8\] \[x^{2.5}-(4200)^{2.5}=2.5*10^8\] \[x^{2.5}=(4200)^{2.5}+2.5*10^8\] \[x=((4200)^{2.5}+2.5*10^8\])^{1/}

OpenStudy (amistre64):

ugh ... firefox crashed at the end of that

OpenStudy (amistre64):

\[x=((4200)^{2.5}+2.5*10^8)^{1/2.5}=4545.ddd\]

OpenStudy (anonymous):

I tried integrals, I must have done it wrong.

OpenStudy (amistre64):

letting x=4545 seems to be fair

OpenStudy (anonymous):

oh, I see what I did wrong. Thank you :D

OpenStudy (amistre64):

yw